# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ IV Pairing from magnons
We turn now to the central question of whether magnon-mediated interactions derived in Sec. III can precipitate superconductivity in a manner compatible with experimental observations in rhombohedral graphene. In what follows, we assume a normal state exhibiting majority Fermi surfaces whose characteristic Fermi momentum greatly exceeds that of the minority Fermi surfaces—as manifested by quantum oscillation measurements in the dominant superconducting regions of spin-orbit-proximitized bilayers and trilayers [37, 38, 40, 39, 41, 42]. Unless specified otherwise, we will postulate generic spin-canting order in the normal state (i.e., with canting angle $\theta_{0}$ not too close to 00 or $\pi/2$).
The pairing problem in this regime poses an interesting puzzle. Naively, one might expect the dominant pairing interaction to emerge directly from the intervalley part of Eq. (33) that scatters a zero-momentum pair of electrons from the majority Fermi surfaces to the minority Fermi surfaces and vice versa [see Fig. 3(a)]. Such processes reflect Cooper pairs interacting through the exchange of a single magnon. The extreme mismatch between the Fermi pocket sizes, however, necessitates that the mediating magnon carries large momentum (comparable to the majority pocket’s Fermi momentum)—thus not utilizing the softness of the Goldstone mode. Numerous issues then arise with this scenario as detailed in Sec. IV.1.
These issues prompt the search for an unconventional pairing mechanism, which we present in Sec. IV.2, that benefits directly from the softness of Goldstone magnons. Our scenario is rooted in the observation that exchange of a single long-wavelength, low-energy magnon flips the spins of a Cooper pair near the Fermi energy and scatters it to high-energy states of the opposite spin [see Fig. 4(a)]. For example, a pair of holes may be transferred from the minority Fermi surfaces to deep within the majority Fermi sea. While these “off-Fermi-surface” scattering events are clearly unimportant for superconductivity at first order in Eq. (33), second-order processes involving two exchanges of low-energy magnons can bring the Cooper pair back to the Fermi energy and thus mediate an effective pairing interaction. We show that, crucially, such a two-magnon interaction exhibits a low-frequency divergence that, owing to its strongly retarded nature, can stabilize superconductivity even in the face of Coulomb repulsion. The superconducting order parameter emerging in this scenario is predicted to be $s$-wave. Moreover, the divergence in the two-magnon interaction requires spin-orbit coupling and vanishes in the $\lambda_{I}\rightarrow 0$ limit, possibly explaining the observed rise in $T_{c}$ with increasing $\lambda_{I}$ in proximitized bilayer samples [37].
Although the two-magnon scenario we propose below is couched in the language of perturbation theory, it is well justified due to its unique singular behavior that qualitatively distinguishes itself from all other magnon-mediated processes—not only the one-magnon process but also all those involving $\geq 3$ magnons. This justification differs from many other perturbation theories that rely on the validity of a large-$N$ expansion.
IV.1 Why the conventional single-boson exchange picture fails
Figure 3: One-magnon exchange processes. (a) Illustration of the momentum-space kinematics and (b) Feynman diagram for one-magnon pair scattering. Such events do not take advantage of low-energy Goldstone modes in the spin-canted phase, and thus do not account for spin-orbit-enabled superconductivity in rhombhedral graphene.
Typically, as in standard phonon-mediated pairing, attraction is generated by exchanging one boson. Our problem also supports a process of this type wherein Cooper pairs exchange one magnon; see Fig. 3(b), which represents a magnon-mediated pair scattering event between majority and minority bands. Note in particular that the spin for each valley flips, as required by the transverse nature of the magnon modes that we have retained. Such a process must in turn transfer a large momentum via the exchanged magnon as highlighted above, leading to various challenges for pairing that we now enumerate.
One clear issue is that the pairing interaction does not directly utilize the ${\boldsymbol{q}}=0$ part of the magnons, and therefore cannot qualitatively distinguish between gapless and gapped branches. The “on-Fermi-surface” Cooper-pair scattering amplitudes arising from exchange of a single magnon of any type are correspondingly non-singular in the low-frequency limit. Consider, for instance, the Goldstone mode labeled $s-$. At small momenta, $|{\boldsymbol{q}}|\lesssim q_{\rm max}$, the $s-$ mode is low-energy and linearly dispersing, and mediates scattering amplitudes calculated in Eq. (37). At larger $|{\boldsymbol{q}}|$, however, $s-$ acquires a quadratic dispersion and bends upward, reaching an energy scale comparable to the Fermi energy at $|{\boldsymbol{q}}|\sim k_{F}$. Repeating the analysis of Sec. III in the latter regime, we find that the typical “on-Fermi-surface” pair scattering amplitude is modified,
$g^{-+,-+}_{K,K^{\prime}}(\nu,|{\boldsymbol{q}}|\sim k_{F})\sim\frac{1}{-i\gamma\nu+Dk_{F}^{2}}.$ (42)
The $k_{F}$ factor invariably preempts any low-frequency divergence in the pairing interaction mediated by one $s-$ magnon; similar conclusions hold for the other branches.
The non-singular nature of the one-magnon pairing interactions already justifies focusing on two-magnon processes—which we show in the next subsection do generate a low-frequency divergence. Nevertheless, it is worth highlighting additional issues with this scenario. As a reminder, to achieve pairing in the presence of strong Coulomb repulsion, usually one invokes the Anderson-Morel picture: pair-breaking effects of Coulomb are suppressed when the Fermi energy $\epsilon_{F}$ greatly exceeds the bandwidth of the pairing interaction (i.e., when the pairing interaction exhibits strong retardation). This regime holds for phonon-mediated attraction, whose bandwidth is set by the Debye frequency that is much lower than $\epsilon_{F}$. However, in the one-magnon exchange scenario, the bandwidth of the pairing interaction is set by the magnon energy which, as noted above, is expected to be comparable to $\epsilon_{F}$. Indeed Eq. (42) is a Lorentzian function of $\nu$ with a bandwidth $\propto\epsilon_{F}$. Thus it is far from obvious that attraction from this subdominant channel could yield superconductivity in the presence of Coulomb repulsion.
Finally, the one-magnon scenario does not naturally explain the observed superconducting phenomenology in graphene multilayers. Only large-momentum magnons participate, whose energy are similar with and without spin-orbit interaction, and regardless of whether the parent normal state exhibits spin-canting order. In sharp contrast, experiments indicate that proximity-induced spin-orbit coupling plays a decisive role in promoting Cooper pairing.
IV.2 Two-magnon exchange scenario
So far we have shown that a traditional scenario where Cooper pairs exchange one magnon does not compellingly explain superconductivity emerging from a spin-canted phase. Can the Goldstone mode then provide a viable pairing glue? In this subsection, we definitively answer this question ‘yes’ and show that, as claimed earlier, the dominant attraction channel arises from two small-momentum magnon exchanges, where the first exchange brings Cooper pairs to high-energy states and the second brings them back to the Fermi level; recall Fig. 4(a).
Let us start by calculating the pairing interaction generated by two-magnon exchanges. Since the bare magnon-mediated interactions from Eq. (33) generate high-energy states at small $q$, one can treat $\mathcal{S}_{I}$ in perturbation theory when focusing on soft-magnon contributions. At second order one obtains an intervalley four-fermion interaction that, after projecting onto the pairing channel, can be expressed as
$\mathcal{S}^{\text{2-mag}}_{\rm pair}=-\sum_{\alpha=\pm}\int_{kq}g^{(2)}_{\Phi}(q)\overline{\Phi}^{\alpha}_{k+q}\Phi^{\alpha}_{k},$ (43)
where we defined the Cooper pair field $\Phi^{\alpha}_{k}=\Psi_{\alpha K^{\prime},-k}\Psi_{\alpha K,k}$, with $\alpha=\pm$ denoting the majority ($+$) or minority (--) band. The coupling function $g_{\Phi}^{(2)}(q)$ turns out to be $\alpha$-independent within our approximations. In the convention specified above, positive values of $g_{\Phi}^{(2)}$ correspond to attractive interactions. We shall see below that $g_{\Phi}^{(2)}$ is indeed positive regardless of the sign of each scattering amplitude because it involves two magnons.
The two-magnon pairing interaction in Eq. (43) can be expressed through irreducible second-order diagrams as shown in Fig. 4; the associated small-$q$ scattering amplitudes $g^{\alpha^{\prime}\alpha,\beta^{\prime}\beta^{\prime}}_{\xi_{1},\xi_{2}}$ are given in Eq. (37). The figure illustrates four types of processes: (a) the ladder diagram and (b) the crossed or “exchange” diagram, that correspond to virtual pair scattering, (c) vertex corrections or “wineglass” diagrams, and (d) bubble diagrams.
Figure 4: Pairing interaction mediated by two-magnon exchanges. Four types of two-magnon processes are given by (a) the ladder diagram, (b) the cross diagram, (c) the vertex corrections and (d) the bubble diagram. Solid lines represent electrons near the Fermi surface that can form Cooper pairs, while dashed lines represent virtual excitations in the intermediate stages; see Fig. 1(b). Wiggly lines represents magnon-mediated interactions of three types. Blue lines ($g^{+-,+-}_{K,K^{\prime}}$ and $g^{-+,-+}_{K,K^{\prime}}$) and red lines ($g^{+-,-+}_{K,K^{\prime}}$ and $g^{-+,+-}_{K,K^{\prime}}$) represent intervalley scatterings, while yellow lines represents intravalley scatterings.
Both vertex corrections and bubble diagrams suffer from the same issue as the first-order diagrams—namely, external electron lines near the Fermi surface enforce a large momentum transfer that does not directly benefit from the long-wavelength Goldstone mode. Vertex corrections thus also exhibit a broad bandwidth and, most importantly, their contribution is non-divergent at low frequency and therefore negligible compared to the other second-order diagrams.
We turn now to the pairing interactions generated by the ladder diagram $g_{\rm l}^{(2)}$ and cross diagram $g_{\rm x}^{(2)}$, which can be expressed as
$g^{(2)}_{\rm l}(q)$ $=\frac{1}{4\bar{m}^{2}}\int_{Q}g^{+-,+-}_{K,K^{\prime}}(Q+q)g^{-+,-+}_{K,K^{\prime}}(-Q),$ (44)
$g^{(2)}_{\rm x}(q)$ $=\frac{1}{4\bar{m}^{2}}\int_{Q}g^{+-,-+}_{K,K^{\prime}}(Q+q)g^{-+,+-}_{K,K^{\prime}}(-Q),$ (45)
where we defined the short-hand $Q\equiv(i\Omega,{\boldsymbol{Q}})$ [59]. Using the small-$q$ asymptotic expressions for $g^{+-,-+}_{K,K^{\prime}}$ and $g^{-+,-+}_{K,K^{\prime}}$ given in Eq. (37), we see that these two diagrams yield identical contributions; together they dominate the two-magnon pairing interaction
$g_{\Phi}^{(2)}(q)=g^{(2)}_{\rm l}(q)+g^{(2)}_{\rm x}(q)$ (46)
$\sim\frac{\cos^{4}\theta_{0}}{8\bar{m}^{2}z_{s}^{2}\kappa^{2}}\int_{Q}\frac{1}{(\Omega+\nu)^{2}+c_{s}^{2}|{\boldsymbol{Q}}+{\boldsymbol{q}}|^{2}}\frac{1}{\Omega^{2}+c_{s}^{2}|{\boldsymbol{Q}}|^{2}}.$
Power-counting provides a shortcut to see that $g_{\Phi}^{(2)}$ diverges as $q\rightarrow 0$. Namely, there are three integrals (one frequency, two momentum components), whereas the denominator of the integrand is quadratic in $\nu$ and ${\boldsymbol{q}}$. Thus the pairing interaction should scale as $\sim 1/\nu$ at small $\nu$ with ${{\boldsymbol{q}}}=0$, and as $\sim 1/|{\boldsymbol{q}}|$ at small ${\boldsymbol{q}}$ with $\nu=0$. We substantiate this argument in Appendix E by evaluating the integrals in the following two extreme regimes:
$g_{\Phi}^{(2)}(q)\sim\begin{cases}\frac{I}{c_{s}|{\boldsymbol{q}}|},&\nu\ll c_{s}|{\boldsymbol{q}}|\\ \frac{I}{\nu},&\nu\gg c_{s}|{\boldsymbol{q}}|\end{cases},~{}~{}I=\frac{V+J}{16z_{s}D}\cos^{2}\theta_{0}.$ (47)
To proceed analytically while preserving the preceding IR divergences, we substitute $g_{\Phi}^{(2)}(q)$ with the expression
$g_{\Phi}^{(2)}(q)\rightarrow\frac{I}{\sqrt{\nu^{2}+c_{s}^{2}|{\boldsymbol{q}}|^{2}}}.$ (48)
As this pairing interaction does not exhibit any sign alternation in momentum space, we focus below on $s$-wave pairing that is expected to dominate. There are two $s$-wave channels to consider: one on the majority Fermi surfaces (+)(+)( + ) and the other on the minority Fermi surfaces ()(-)( - ). The effective $s$-wave interactions in these two channels are simply given by the average of $g_{\Phi}^{(2)}(q)$ over the corresponding Fermi surfaces:
$\overline{g_{\Phi,\pm}^{(2)}}(\nu)$ $=\int\frac{dq_{\parallel}}{2\pi k_{F}^{\pm}}g_{\Phi}^{(2)}(\nu,(q_{\parallel},q_{\perp}=0)).$ (49)
Here $q_{\parallel}$ and $q_{\perp}$ denote momentum components parallel and normal to the Fermi surface exhibiting Fermi momentum $k_{F}^{\alpha}$ in band $\alpha$. Upon working out this integral we find a logarithmic divergence in the low-frequency limit,
$\overline{g_{\Phi,\pm}^{(2)}}(\nu)\sim\frac{\cos\theta_{0}}{64\pi\bar{m}k_{F}^{\pm}}\sqrt{\frac{V+J}{z_{s}D^{3}}}\ln\left(\frac{c_{s}q_{\rm max}}{|\nu|}\right),$ (50)
with $q_{\rm max}=\frac{1}{V}\sqrt{\frac{J}{D}}\cos\theta_{0}$ the momentum cutoff below which the dispersion of the $s-$ magnon mode is linear. This interaction is attractive and therefore catalyzes $s$-wave pairing mediated by two-magnon exchange. Importantly, the logarithmic divergence at low frequency renders the interaction narrow-band, and therefore pairing resulting from this mechanism more readily survives bare Coulomb repulsion through the Anderson-Morel mechanism.
The $\cos\theta_{0}$ factor in Eq. (50) translates into a linear dependence on $\lambda_{I}$ using the saddle-point condition Eq. (13), thus linking spin-orbit coupling with enhanced pairing. In the $\theta_{0}\rightarrow 0$ limit where spin-canting order disappears in favor of spin-valley locking, we have shown in Sec. III.2 that the additional gapless (critical) mode at the transition yields intervalley magnon interactions with the same frequency and momentum structure as in the generic canting case, albeit with different amplitudes. In this limit only the ladder diagrams in Fig. 4(b) contribute, ultimately leading to a pairing interaction identical to Eq. (50) but with $\cos\theta_{0}\rightarrow 1$. Our mechanism thus predicts a saturation of the pairing strength when increasing $\lambda_{I}$ towards the canting transition. However, when the spin-valley locked phase is nucleated, we expect superconductivity to quickly evaporate as all collective modes becomes gapped—see also Fig. 2.
Within the weak-coupling BCS framework, the $s$-wave pairing gap $\Delta_{\pm}$ for the majority ($+$) and minority (--) bands is determined by the gap equations
$\Delta_{\pm}(\omega)=\mathcal{N}_{\pm}\int_{\nu}\overline{g_{\Phi,\pm}^{(2)}}(\omega-\nu)\frac{\Delta_{\pm}(\nu)}{\sqrt{\nu^{2}+\Delta_{\pm}^{2}(\nu)}},$ (51)
with $\mathcal{N}_{\pm}$ the density of states at the corresponding Fermi surfaces. Here, we only account for the leading contribution from two-magnon interactions; moreover, we have neglected self-energy corrections (which we discuss in Sec. V). The two gaps $\Delta_{\pm}$ decouple at this level, though they will mix by subleading interactions (e.g., one-magnon exchange) that provide residual Josephson coupling between the bands. We stress that $\overline{g_{\Phi,\pm}^{(2)}}$ is inversely proportional to the Fermi surface radius; recall Eq. (50). This unusual feature arises because the zero-frequency pairing interaction is strongly peaked near momentum transfer ${\boldsymbol{q}}=0$—in contrast to, e.g., acoustic phonons for which $g_{\rm ph}(\nu=0,{\boldsymbol{q}})$ is nearly momentum-independent. Assuming the density of states $\mathcal{N}_{\pm}$ are comparable, stronger pairing is correspondingly expected for the minority carriers. This prediction is in harmony with observations in spin-orbit-proximitized graphene multilayers, where superconductivity appears in parameter regimes featuring minority pockets; for further discussion on this important point see Sec. VI.
We have now seen that the one-magnon scenario faces issues that two-magnon processes remedy. A crucial question remains, however: Can higher-order magnon-exchange events produce even stronger contributions, potentially dooming the entire perturbative framework? Fortunately, higher-order diagrams do not exhibit a low-frequency divergence and are thus negligible compared to those that involve two magnons. This result can be seen by power counting: Introducing one more magnon line into the diagrams in Fig. 4 yields two additional powers of momentum and frequency in the denominator of the integrand—yet also introduces extra integrals along three dimensions (one frequency, two momentum components). Higher-order diagrams are therefore less divergent compared to those arising at second order. Given that $\overline{g_{\Phi}^{(2)}}(\nu)$ is merely logarithmically divergent, we conclude that the momentum-averaged pairing interaction arising from higher-order diagrams is indeed non-divergent.
IV.3 Suppression of pairing interaction in the absence of spin-orbit coupling
In the previous subsection we considered generic canting angles as well as the special limit $\theta_{0}\rightarrow 0$. Here we focus on the $\theta_{0}\rightarrow\pi/2$ limit, corresponding to a vanishing spin-orbit interaction $\lambda_{I}\rightarrow 0$, which may be relevant for experiments without a neighboring TMD. In this limit, the leading contribution to the interaction in Eq. (50) vanishes due to the $\cos\theta_{0}$ factor. The physical interpretation of this behavior can be traced back to the disappearance of linearly dispersing Goldstone magnons; the magnon dispersion becomes quadratic as the microscopic Hamiltonian preserves SU(2) rather than U(1) symmetry.
To understand the fate of pairing in the $\theta_{0}=\pi/2$, spin-orbit-free limit, we calculate the two-magnon interaction generated by the quadratically dispersing modes at small ${\boldsymbol{q}}$ [recall Eq. (41)]. Only the crossed diagrams in Fig. 4(b) contribute, due to symmetry considerations described above that kill vertices that change the total spin projection along the in-plane polarization axis. The pairing interaction thus reads
$g^{(2)}_{\Phi,\frac{\pi}{2}}(q)\sim\frac{1}{4\bar{m}^{2}}\int_{Q}\frac{1}{-i\gamma(\Omega+\nu)+D({\boldsymbol{Q}}+{\boldsymbol{q}})^{2}}\frac{1}{i\gamma\Omega+D{\boldsymbol{Q}}^{2}}.$ (52)
The above integral can be computed exactly; see Appendix E.2. Focusing on $\nu=0$ (i.e., magnon-mediated interactions resonant on the Fermi surface), we obtain
$g^{(2)}_{\Phi,\frac{\pi}{2}}(\nu=0,{\boldsymbol{q}})\sim\frac{V+J}{8\pi D\bar{m}}\ln\frac{2\Lambda}{|{\boldsymbol{q}}|},$ (53)
where $\Lambda$ is a momentum cut-off related to the finite bandwidth of the quadratically dispersing magnon modes.
Contrary to the generic spin-canting result in Eq. (48), this zero-frequency interaction is only logarithmically divergent at small momentum. Consequently, its momentum-averaged value is finite and subdominant relative to the $\lambda_{I}\neq 0$ case. That is, spin-orbit interaction qualitatively enhances $s$-wave superconductivity mediated by two-magnon exchange—also in harmony with rhombohedral graphene experiments.
V Does Pairing survive self-energy?
In Sec. IV.2 we established that the momentum-averaged pairing interaction arising from the two-magnon scenario is attractive and log-divergent at low frequency, yielding $s$-wave superconductivity within a weak-coupling BCS framework (see Eq. 51). Caution is warranted, however, as the presence of a singular pairing interaction might also lead to a singular imaginary part of the electron self-energy $\Sigma(i\nu)$. Furthermore, $\Sigma(i\nu)$ affects the coherence of fermions and blurs the Fermi surface—thus generally competing with pairing, as the latter relies on the energy gain from gapping out a sharp Fermi surface. A similar situation appears in quantum critical pairing scenarios [60, 61, 62, 63, 64, 65, 66, 67, 68, 68, 69], where a quantum critical mode mediates a singular pairing interaction as well as a singular imaginary part of the self-energy. In this section we address two questions: Is the self-energy actually singular in our problem? And if so, does the tendency towards pairing or towards non-Fermi liquid behavior win out?
Figure 5: Self-energy from magnon-mediated interactions. (a) Two types of intravalley scattering processes appearing in the self-energy. (b) Self-energy contributions arise from a one-magnon effective interaction $g_{\Sigma}^{(1)}$, a two-magnon interaction $g_{\Sigma}^{(2)}$, and those from higher orders (denoted by the ellipsis). The one-magnon piece $g^{(1)}_{\Sigma}$ simply corresponds to the intravalley scattering amplitude $g^{+-,-+}_{K,K}$. The two-magnon piece $g^{(2)}_{\Sigma}$ (double wiggly black lines) is given by the two diagrams in panel (c). These diagrams are similar to those in the two-magnon pairing interaction—recall Fig. 4—except that only intravalley scattering processes are involved.
Unlike in problems studied in the literature such as Refs. 70, where the divergent pairing interaction directly generates a singular self-energy, in our problem the self-energy originates from a series of processes distinct from the pairing interaction. Figure 5 illustrates the self-energy-generating diagrams at each order $g_{\Sigma,\alpha}^{(n)}$ ($n=1,2,...$), which are structurally similar to those of the pairing interaction at order $g_{\Phi}^{(n)}$. However, the magnon lines (wiggly lines) in the pairing and self-energy correspond to different scattering types. In the pairing interaction, wiggly lines represent intervalley processes whereas in the self-energy, wiggly lines represent intravalley processes. For generic $\theta_{0}$ values away from 0 and $\pi/2$, this distinction nevertheless does not affect the leading divergence of the self-energy: As shown in Sec. III.1 the asymptotic expressions for intervalley $g_{K,K^{\prime}}$ and intravalley $g_{K,K}$ amplitudes are identical [Eq. (37)].
We thus conclude that the interactions $g_{\Sigma,\alpha}^{(n)}$ that govern the self-energies are asymptotically identical to the pairing interactions $g_{\Phi,\alpha}^{(n)}$. The leading contribution also occurs at the two-magnon level [see Fig. 5(c)], with $g_{\Sigma,\alpha}^{(2)}$ diverging logarithmically at low frequency. A straightforward calculation shows that the self-energies $\Sigma_{\alpha}$ ($\alpha=\pm$) feature the following leading singularity,
[…]
where $\omega_{*}$ is a frequency cutoff and $\overline{g_{\Sigma,\alpha}^{(2)}}$ denotes an average of the corresponding amplitude over the $\alpha$ Fermi surface. In contrast, the first-order diagram in Fig. 5(b) does not give a singular self-energy contribution because the interaction scatters an electron between the $+$ and -- Fermi surfaces, which requires a large momentum transfer.
Interestingly, the situation of ${\rm\,Im\,}\Sigma(i\omega)\sim\omega\ln|\omega|$ described in Eq. (54) shows that the normal (non-superconducting) state realizes a marginal Fermi liquid[71]. This situation resembles the rather different problem of conventional boson-mediated pairing in three dimensions. Reference 72 in particular established that exchanging one boson similarly mediates a log-divergent pairing interaction as well as a self-energy with $\omega\ln|\omega|$ singularity—though $T_{c}$ for pairing is still nonvanishing. We therefore conclude that pairing also occurs in our two-dimensional, two-magnon-exchange scenario despite the singular self-energy. We note that, since $g_{\Sigma}^{(n)}$ is asymptotically identical to $g_{\Phi}^{(n)}$, the fate of the self-energy generated at third-order or higher is similar to that of the high-order irreducible pairing interaction—such corrections are also non-divergent and therefore negligible compared to the two-magnon ones.
VI Discussion
Our theory admits several experimentally testable predictions. We start by discussing the dependence of superconductivity on Ising spin-orbit coupling strength $\lambda_{I}$, which naturally accommodates two seemingly contradictory behaviors:

### v3

@@ IV Pairing from magnons
Can the spin-flipping pair scattering process described above give rise to superconductivity? In this Section, we study this possibility and estimate the corresponding superconducting $T_{c}$.
Due to the spin-flipping nature of the magnon-mediated interaction in Eqs. (38), the pairing process couples the superconducting gaps in the two (majority and minority) spin flavors, which we denote by $\Delta_{+}$ and $\Delta_{-}$, respectively. The linearized gap equation is given by
$\Delta_{+}(k)=-\sum_{k^{\prime}}g^{-+,-+}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{-}(k^{\prime})}{\omega_{n}^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},-}^{2}},$ (40)
$\Delta_{-}(k)=-\sum_{k^{\prime}}g^{+-,+-}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{+}(k^{\prime})}{\omega_{n}^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},+}^{2}},$ (41)
where $k=(\omega,{\boldsymbol{k}})$. Below we focus on the momentum-independent, $s$-wave pairing channel (see however footnote 1), justified by the smooth momentum dependence of the pairing interaction. Due to the interband nature of the pairing, the coupled equations (40) admit a momentum-independent solution regardless of the sign of $g^{-+,-+}_{K,K^{\prime}}$. In our case, $g^{-+,-+}_{K,K^{\prime}}>0$ represents a repulsive interaction, and pairing can thus occur in a channel with opposite signs of $\Delta_{+}$ and $\Delta_{-}$, similar to $s_{+-}$ pairing in iron-based superconductors (see, e.g., Ref. 62).
To proceed analytically, we apply the trick conventionally used in solving the BCS problem: replacing the full pairing interaction with a frequency-independent version that exhibits a hard cut-off at the characteristic magnon energy $\Omega_{0}$ (analogous to the Debye energy in conventional phonon-mediated pairing). For simplicity, we focus on the small canting angle regime so that, up to first order in $J/V$, the two magnon velocities are similar, $c_{s}\sim c_{a}\sim c=\sqrt{\frac{JD}{V\kappa}}$, and $z_{s}\sim z_{a}\sim z=\frac{V}{2J}$. Here the characteristic magnon energy is estimated as $\Omega_{0}=ck_{0}$, where $k_{0}$ is a characteristic scale for the momentum transfer ${\boldsymbol{k}}-{\boldsymbol{k}}^{\prime}$ in the spin-flipping scattering process. This scale is upper-bounded by $2k_{F}^{+}$, with $k_{F}^{+}$ the Fermi momentum for the majority Fermi surface. However, in rhombohedral graphene the momentum-space separation between majority and minority pockets can be several times smaller than $k_{F}^{+}$, due in particular to trigonal warping (e.g., Ref. 53).
With these considerations in mind, we apply the following replacement:
$\frac{1}{(\omega_{n}-\omega^{\prime}_{n})^{2}+c^{2}(k-k^{\prime})^{2}}$ $\rightarrow\frac{\Theta(\Omega_{0}-|\omega_{n}^{\prime}|)\Theta(\Omega_{0}-|\omega_{n}|)}{\Omega_{0}^{2}},$
$\Delta_{\pm}(\omega_{n},k)$ $\rightarrow\Delta_{\pm}^{(0)}\Theta(\Omega_{0}-|\omega_{n}|),$ (42)
which reduces the linearized gap equation to an analytically soluble form,
$\Delta_{+}^{(0)}=-g_{\rm eff}\nu_{-}\Delta_{-}^{(0)}\log\frac{\Omega_{0}}{T_{c}},$ (43)
$\Delta_{-}^{(0)}=-g_{\rm eff}\nu_{+}\Delta_{+}^{(0)}\log\frac{\Omega_{0}}{T_{c}},$ (44)
with an effective pairing strength
$g_{\rm eff}\approx\frac{\cos^{2}\theta_{0}}{\frac{V}{2J}\kappa\Omega_{0}^{2}}=\frac{2\cos^{2}\theta_{0}}{Dk_{0}^{2}}.$ (45)
Defining an effective density of state $\nu_{\rm eff}=\sqrt{\nu_{+}\nu_{-}}$, the superconducting $T_{c}$ solved from Eqs. (43) and (44) reads
$T_{c}=\Omega_{0}\exp{\left(-\frac{1}{g_{\rm eff}\nu_{\rm eff}}\right)}.$ (46)
This result enables a crude $T_{c}$ estimate. The dimensionless coupling constant $g_{\rm eff}\nu_{\rm eff}$ is independent of the ratio $J/V$ and only involves band structure quantities. Therefore, it is expected to be of $\mathcal{O}(1)$ from dimensional analysis, but will depend on the fermiology and density of states of a specific system. The magnon energy $\Omega_{0}$ is proportional to a small dimensionless ratio $\sqrt{J/V}$ times the Fermi energy (again assuming that $\bar{m}\sim\epsilon_{F})$. In realistic systems, we expect $V/J\sim O(10)$. Therefore, the superconducting $T_{c}$ predicted in our scenario can reach a substantial fraction of $\epsilon_{F}$. Given that we neglected the competing effect of repulsive Coulomb interactions, this estimate should be understood as an upper bound on $T_{c}$. Indeed, in lightly doped rhombohedral bilayers and trilayers (where $\epsilon_{F}\sim 1-10$ meV), the reported critical temperatures [40, 42, 41, 44, 43, 45] for the spin-orbit enabled superconductors are only in the range $k_{B}T_{c}\sim 10-50\mu$eV.
V Discussion
Our theory admits several experimentally testable predictions. Perhaps the most significant and unique one is the dependence of superconductivity on Ising spin-orbit coupling strength $\lambda_{I}$. Qualitatively, the pairing interaction in our theory vanishes by symmetry when $\lambda_{I}$ is switched off, which can elegantly explain the observed superconducting phases in TMD-proximitized rhombohedral bilayer/trilayer systems that rely on induced spin-orbit coupling [40, 41, 42, 43, 44, 45]. Our theory further naturally accommodates two seemingly contradictory dependences on spin-orbit coupling, which cannot be easily understood in a conventional (e.g., phonon-mediated) pairing mechanism: